OpenStudy (nikki17_guess):

Credit card A offers an APR of 25.68%, compounded monthly, while credit card B offers an APR of 25.32%, compounded daily. All else being equal, which card offers the better deal for the consumer? A.Credit card B, because its effective interest rate is about 0.13% greater than that of credit card A. B. Credit card B, because its effective interest rate is about 0.13% less than that of credit card A. C. Credit card A, because its effective interest rate is about 0.13% less than that of credit card B. D. Credit card A, because its effective interest rate is about 0.13% greater than tha

ganeshie8 (ganeshie8):

Work the quantities \[\left(1+\dfrac{0.2568}{12}\right)^{12}\] and \[\left(1+\dfrac{0.2532}{365}\right)^{365}\]

OpenStudy (nikki17_guess):

still don't get it